OpenStudy (anonymous):

What is an absolute value equation with a vertex at (-3,4)

OpenStudy (jmark):

Given vertex at (-3,4) =(h,k) we know y=m|x-h|+k is the parent equation. m=slope, h=where the vertex is shifted (left or right) from the origin, in this case it is shifted 3 units to the left k=where the vertex is shifted (up or down) from the origin, in this case it is shifted 4 units up =>absolute value equation =>y=m|x+3|+4 let (x,y)=(0,0) =>0=m|0+3|+4 =>m=-4/3 hence absolute value equation =>y= -4/3|x+3|+4 See also: CBSE Sample Papers http://goo.gl/Z5tWIh